can you pleaseeeee help me..
My problem says to state the domain for the function f (x)= x-2 divided by x+2
i dont understand how to do this at all.. if you could help i would greatly appreciate it

Answer by stanbon(75887)

state the domain for the function f (x)= x-2 divided by x+2
-----------
f(x)=(x-2)/(x+2)
x can take on any value except one that would make the denominator
be zero.
If x+2=0, the x=-2
So the domain is "all Real numbers except x=-2".
Cheers,
Stan H.